# Differentially Private Optimization with Sparse Gradients

### v1

@@ Abstract
Motivated by applications of large embedding models, we study differentially private (DP) optimization problems under sparsity of individual gradients. We start with new near-optimal bounds for the classic mean estimation problem but with sparse data, improving upon existing algorithms particularly for the high-dimensional regime. Building on this, we obtain pure- and approximate-DP algorithms with almost optimal rates for stochastic convex optimization with sparse gradients; the former represents the first nearly dimension-independent rates for this problem. Finally, we study the approximation of stationary points for the empirical loss in approximate-DP optimization and obtain rates that depend on sparsity instead of dimension, modulo polylogarithmic factors.
@@ 1 Introduction
Consider neural network models whose inputs have categorical features with large vocabularies. A typical way for dealing with such features is to use embedding tables; namely, for a feature that takes $K$ distinct values, we create trainable parameters $w_{1},\ldots,w_{K}\in\mathbb{R}^{k}$, and use $w_{a}$ as input to the neural network when the corresponding input feature is $a$. A natural outcome of considering such models is that the per-example gradients are guaranteed to be sparse; when the input feature is $a$, then only the gradient with respect to $w_{a}$ is non-zero. Given the prevalence of sparse gradients in practical deep learning applications, GPUs/TPUs that are optimized to leverage gradient sparsity are commercially offered and widely used in industry [22, 38, 1, 25]. To leverage precisely this sparsity of gradients, recent practical work has considered DP stochastic optimization with sparse gradients for large embedding models for different applications including recommendation systems, natural language processing, and ads modeling [41, 21].
Despite its relevance and promising empirical results, there is limited understanding of the theoretical limits of DP learning under gradient sparsity. In this work, we propose a theoretical framework for studying this problem, together with nearly tight bounds for various settings.
@@ 1.1 Our Results
Our results unearth three regimes of accuracy rates for the above setting: (i) the small dataset size regime where the optimal rate is constant, (ii) the large dataset size where the optimal rates are polynomial in the dimension and (iii) an intermediate dataset size regime where a high-dimensional rate can be obtained [^1] We will generally refer to high-dimensional or nearly dimension-independent rates indistinguishably, meaning more precisely that the rates scale poly-logarithmically with the dimension. (this roughly corresponds to the square root of the classical rate, where $d$ is replaced by $s\log d$; see Table 1 and Table 2, for the precise rates). These results imply in particular that even for high-dimensional models, this problem is tractable under gradient sparsity. Interestingly, without sparsity, these poly-logarithmic rates are unattainable in the light of known lower bounds [6].
@@
Convex $(\varepsilon,\delta)$-DP ERM ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{(s\ln(d)\ln(1/\delta))^{1/4}}{\sqrt{\varepsilon n}}}$ (Thm. 5.1) $\frac{\sqrt{d\ln(1/\delta)}}{\varepsilon n}$
@@
Convex $\varepsilon$-DP ERM ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Big{(}\frac{s\ln(d)}{\varepsilon n}\Big{)}^{1/3}}$    (Thm. 5.1, 5.5) $\frac{d}{\varepsilon n}$
@@
Noncvx. $(\varepsilon,\delta)$-DP Emp. Grad. Norm ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{(s\ln(d/s)\ln^{3}(1/\delta))^{1/8}}{(\varepsilon n)^{1/4}}}$ (Thm. 6.8) $\Big{(}\frac{\sqrt{d\ln(1/\delta)}}{\varepsilon n}\Big{)}^{2/3}$
@@
Table 2: Rates for high-dimensional DP optimization with sparse gradients, compared to best-existing upper bounds in the non-sparse case. Bounds stated for constant success probability. Function parameters and polylog$(n)$ factors omitted. Improvements ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{highlighted}}$.
@@
In Section 3, we start with the fundamental task of $\ell_{2}$-mean estimation of a dataset with sparse data (which reduces to ERM with sparse linear losses [6]). Here, we obtain new upper bounds (see Table 1). These rates are obtained by adapting the projection mechanism [31], with a convex relaxation that makes our algorithms efficient. Note that for pure-DP, even our low-dimensional rate of $\sqrt{sd}/(\varepsilon n)$ can be substantially smaller than the dense pure-DP rate of $d/(\varepsilon n)$ [6], whenever $s\ll d$. For approximate-DP we also obtain a sharper upper bound by solving an $\ell_{1}$-regression problem of a noisy projection of the empirical mean over a random subspace. Its analysis combines ideas from compressed sensing [40] with sparse approximation via the Approximate Carathéodory Theorem [32].
@@
In Section 4, we prove lower bounds that show the near-optimality of our algorithms. For pure-DP, we obtain a new lower bound of $\Omega(s\log(d/s)/(n\varepsilon))$, which is based on a packing of sparse vectors. While this lower bound looks weaker than the standard $\Omega(d/(n\varepsilon))$ lower bound based on dense packing [23, 6], we design a novel bootstrapping via a block diagonal construction where each block contains a sparse lower bound as above. This, together with a padding argument [6], yields lower bounds for the three regimes of interest. For approximate-DP, we also use the block diagonal bootstrapping, where this time the blocks use classical fingerprinting codes in dimension $s$ [6, 34]. Our lower bounds, however, have a gap of $\ln(d/s)^{1/4}$ in the high-dimensional regime; we conjecture that the aforementioned compressed sensing-based upper bound is tight.
In Section 5, we study DP-SO and DP-ERM in the convex case. For approximate-DP, our upper bounds are obtained by regularized output perturbation with an $\ell_{\infty}$ projection post-processing step. While this projection step is rather unusual, its role is clear from the analysis: it leverages the $\ell_{\infty}$ bounds of noise addition, which in conjunction with convexity provides an error guarantee that also leverages the gradient sparsity. These algorithms are nearly-optimal, by the aforementioned lower bounds. For pure-DP, we provide two different algorithms: one based on regularized output perturbation (which is efficient, but it requires an additional smoothness assumption), and one based on the exponential mechanism [29] run over a net of suitably sparse vectors (which is inefficient, but does not require smoothness). Neither of these algorithms matches the lower bound for mean estimation (the gap in the exponent of the rate is of $1/6$), but they attain the first nearly dimension-independent rates for this problem. Finally, we prove that our DP-ERM bounds transfer to DP-SO guarantees, by virtue of the stability of regularized ERM.
@@
In Section 6 we study the approximation of stationary points in DP-ERM. We propose the use of stochastic gradient (SGD) with a mean estimation gradient oracle based on the results in Section 3. The main challenge here is the presence of bias in mean estimation, which dramatically deteriorates the rates of SGD. In view of this, we propose a bias-reduction method inspired by the simulation literature [9]. This bias-reduction technique uses a random batch size in an exponentially increasing schedule and a telescopic estimator of the gradient which—used in conjunction with our DP mean estimation methods—provides a stochastic first-order oracle that attains bias similar to the one of a full-batch algorithm, with moderately bounded variance. Note that using the full-batch in this case would lead to polynomially weaker rates for stationarity; in turn, our method leverages the batch randomization to conduct a more careful privacy accounting based on subsampling and the fully-adaptive properties of DP [39]. The resulting rate (see Table 1) is approximately the square-root of the rate of noisy SGD in the dense setting where the dimension is replaced by $s\ln(d/s)$, i.e., the desired high-dimensional rate. A plain application of these ideas leads to only constant success probability results, and hence in Section 7 we propose a private confidence boosting procedure to exponentially amplify the success probability.
@@ 1.2 Related Work
DP optimization is an extensively studied topic for over a decade (see Bassily et al. [6, 7], Feldman et al. [20], and the references therein). In this field, some works have highlighted the role of feature sparsity (e.g., using sparsity-promoting $\ell_{1}$-ball constraints) in near-dimension independent excess-risk rates for DP optimization, both for ERM and SCO [24, 35, 36, 4, 8, 15, 14]. These settings are unrelated to ours, as sparse predictors are typically related to dense gradients.
Another proposed assumption to mitigate the impact of dimension in DP learning is that gradients lie in a small dimensional subspace [42, 26]. This is an arguably strong assumption that is not easy to verify, and therefore requires systematic empirical validation. Moreover, the determination of the subspace where gradients need to be projected requires access to nontrivial amounts of public data, which depending on the application may be restrictive.
Substantially less studied is the role of gradient sparsity. Closely related to our work, Zhang et al. [41] studied DP-ERM under gradient sparsity, with some stronger assumptions. Aside from an additional $\ell_{\infty}$ assumption on the individual gradients, the following partitioning sparsity assumption is imposed: the dataset $S$ can be uniformly partitioned into subsets $S_{1},\ldots,S_{m}$, where each of these sets satisfies a uniform bound on gradient sparsity: for all $k\in[m]$ and $x\in{\mathcal{X}}$, $\big{\|}\sum_{z\in S_{k}}\nabla f(x,z)\big{\|}_{0}\leq c_{1}$. The work shows poly-logarithmic in the dimension rates, for both convex and nonconvex settings: for convex objectives the rate is $\frac{\mbox{\footnotesize polylog}(d)}{\varepsilon\sqrt{n}}$, and in the nonconvex case is $\frac{\mbox{\footnotesize polylog}(d)}{\sqrt{\varepsilon}n^{1/4}}$. Our results only assume individual gradient sparsity, so on top of being more general, they are also faster and provably nearly optimal in the convex case.
Another relevant work is Ghazi et al. [21], which studies the computational and utility benefits for DP with sparse gradients in neural networks with embedding tables. With the caveat that variable selection on stochastic gradients is performed at the level of contributing buckets (i.e. rows of the embedding table), rather than on gradient coordinates, this work shows substantial improvements on computational efficiency (as the bucket selection leads to a more efficient batch gradient computation) and also on the resulting utility.
@@ 1.3 Future Directions
First, we conjecture that for approximate-DP mean estimation—similarly to the pure-DP case—a lower bound $\Omega\big{(}\sqrt{s\log(d/s)\ln(1/\delta)}/[n\varepsilon]\big{)}$ should exist; such construction could be bootstrapped with a block-diagonal dataset (more precisely, Lemma 4.3), proving the optimality of our algorithms. Second, for pure-DP, we believe there should exist an algorithm attaining rates analog to those for mean estimation. It appears however that most of the variants of the output perturbation approach (including the use of phasing [20, 4, 5]) cannot attain such rates.
We believe that our bias-reduction approach may find further use in DP. For example, [28, 30] have identified strong limitations from bias in mean estimation. While similar lower bounds may hold for sparse estimation, bias-reduction allows us to amortize this error within an iterative method, preventing their error accumulation. Further exploring this phenomenon may help circumventing the presence of bias for other settings in DP.
@@ 3 Upper Bounds for DP Mean Estimation with Sparse Data
In this section we study the problem of DP mean estimation with sparse data. Our main result is that the projection mechanism [31] is nearly optimal, both for pure- and approximate-DP. In our case, we interpret the marginals on each of the $d$-dimensions as the queries of interest: this way, the $\ell_{2}$-error on private query answers corresponds exactly to the $\ell_{2}$-norm estimation error. A key difference to the approach in Nikolov et al. [31] and related works is that we project the noisy answers onto the set ${\cal K}:={\cal B}_{1}^{d}(0,L\sqrt{s})$, which is a (coarse) convex relaxation of $\mathrm{conv}({\cal S}_{s}^{d})$. This is crucial to make our algorithm efficiently implementable.
@@
return $\hat{z}=\operatorname*{\mathrm{argmin}}\limits_{z\in{\cal K}}\|z-\tilde{z}\|_{2}$, where ${\cal K}:={\cal B}_{1}^{d}(0,L\sqrt{s})$
@@ 3.1 Sharper Approximate DP Mean Estimation via Compressed Sensing
We propose a sharper algorithm for approximate-DP mean estimation based on compressed sensing [40]. The precise improvements relate to reducing the $\log(d)$ factor to $\log(d/s)$, and a faster rate dependence on the confidence $\beta$. The idea is that for sufficiently high dimensions a small number of random measurements suffices to estimate a noisy and approximately sparse signal. These properties follow from existing results in compressed sensing, which provide guarantees based on the $\ell_{2}$-norm of the noise, and the best sparse approximation in the $\ell_{2}$-norm (known as $\ell_{2}\text{-}\ell_{2}$-stable and noisy recovery) [40]. We will exploit such robustness in two ways: regarding the noise robustness, this property is used in order to perturb our measurements, which will certify the privacy; on the other hand, the approximate recovery property is used to find a sparser approximation of our empirical mean. As the approximation is only used for analysis, we can resort on the Approximate Caratheodory Theorem to certify the existence of a sparse vector whose sparsity increases more moderately with $n$ than the empirical average [32] .
@@
Vector $\bar{z}(S)=\frac{1}{n}\sum_{i\in[n]}z_{i}\in\mathbb{R}^{d}$ from dataset $S\in({\cal S}_{s,d})^{n}$; privacy parameters $\varepsilon,\delta>0$
@@
Let $m\eqsim n\varepsilon\sqrt{s\ln(d/s)}/\sqrt{\ln(1/\delta)}$. If $6\exp\{-cm\}\leq\delta<\frac{s\ln(d/s)}{m^{2}}$ (where $c>0$ is a constant) and $0<\varepsilon\leq 1$, then Algorithm 2 is $(\varepsilon,\delta)$-DP, and with probability $1-\delta/2-\beta$,
@@
We now proceed to the accuracy guarantee. By Wojtaszczyk [40, Theorem 3.6 (b)], under the same event as stated above (which has probability $21-\delta/$ 2) we have

### v2

@@ Abstract
Motivated by applications of large embedding models, we study differentially private (DP) optimization problems under sparsity of individual gradients. We start with new near-optimal bounds for the classic mean estimation problem but with sparse data, improving upon existing algorithms particularly for the high-dimensional regime. The corresponding lower bounds are based on a novel block-diagonal construction used in combination with existing DP mean estimation lower bounds. Next, we obtain pure- and approximate-DP algorithms with almost optimal rates for stochastic convex optimization with sparse gradients; the former represents the first nearly dimension-independent rates for this problem. Furthermore, by introducing novel analyses of bias-reduction in mean estimation and randomly-stopped biased SGD we obtain nearly dimension independent rates for near stationary points for the empirical risk in nonconvex settings under approximate-DP.
@@ 1 Introduction
Consider neural networks whose inputs have categorical features with large vocabularies. These features can be modeled using embedding tables; namely, for a feature that takes $K$ distinct values, we create trainable parameters $w_{1},\ldots,w_{K}\in\mathbb{R}^{k}$, and use $w_{a}$ as input to the neural network when the corresponding input feature is $a$. A natural outcome of such models is that the per-example gradients are guaranteed to be sparse; when the input feature is $a$, then only the gradient with respect to $w_{a}$ is non-zero. Given the prevalence of sparse gradients in practical deep learning applications, GPUs/TPUs that are optimized to leverage gradient sparsity are commercially offered and widely used in industry [22, 43, 1, 25]. To leverage gradient sparsity, recent practical work has considered DP stochastic optimization with sparse gradients for large embedding models for different applications including recommendation systems, natural language processing, and ads modeling [46, 21].
Despite its relevance and promising empirical results, there is limited understanding of the theoretical limits of DP learning under gradient sparsity. This gap motivates our work.
@@ 1.1 Our Results
Our results unearth three regimes of accuracy rates for the above setting: (i) the small dataset size regime where the optimal rate is constant, (ii) the large dataset size where the optimal rates are polynomial in the dimension and (iii) an intermediate dataset size regime characterized by a new high-dimensional rate [^1] We will generally refer to high-dimensional or nearly dimension-independent rates indistinguishably, meaning more precisely that the rates scale poly-logarithmically with the dimension. (see Table 1 and Table 2, for precise rates). These results imply in particular that even for high-dimensional models, this problem is tractable under gradient sparsity. Without sparsity, these poly-logarithmic rates are unattainable in the light of known lower bounds [6].
@@
[…]
@@
Table 2: Rates for DP optimization with sparse gradients, compared to best-existing upper bounds in the non-sparse case. Bounds stated for constant success probability. Function parameters and polylog$(n)$ factors omitted. Above ${\cal R}_{\varepsilon,\delta}={\sqrt{d\ln(1/\delta)}}/{[\varepsilon n]}$ and ${\cal R}_{\varepsilon}={d}/{[\varepsilon n]}$. Improvements ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{highlighted}}$.
@@
In Section 3, we start with the fundamental task of $\ell_{2}$-mean estimation with sparse data (which reduces to ERM with sparse linear losses [6]). Here, we obtain new upper bounds (see Table 1). These rates are obtained by adapting the projection mechanism [35], with a convex relaxation that makes our algorithms efficient. Note that for pure-DP, even our large dataset rate of $\sqrt{sd}/(\varepsilon n)$ can be substantially smaller than the dense pure-DP rate of $d/(\varepsilon n)$ [6], whenever $s\ll d$. For approximate-DP we also obtain a sharper upper bound by solving an $\ell_{1}$-regression problem of a noisy projection of the empirical mean over a random subspace. Its analysis combines ideas from compressed sensing [45] with sparse approximation via the Approximate Carathéodory Theorem [36].
@@
In Section 4, we prove lower bounds that show the near-optimality of our algorithms. For pure-DP, we obtain a new lower bound of $\Omega(s\log(d/s)/(n\varepsilon))$, which is based on a packing of sparse vectors. While this lower bound looks weaker than the standard $\Omega(d/(n\varepsilon))$ lower bound based on dense packings [23, 6], we design a novel bootstrapping via a block diagonal construction where each block contains a sparse lower bound as above. This, together with a padding argument [6], yields lower bounds for the three regimes of interest. For approximate-DP, we also use the block diagonal bootstrapping, where this time the blocks use classical fingerprinting codes in dimension $s$ [6, 39]. Our approximate-DP lower bounds, however, have a gap of $\ln(d/s)^{1/4}$ in the high-dimensional regime; we conjecture that the aforementioned compressed sensing-based upper bound is tight.
@@
In Section 5, we study DP-ERM with sparse gradients, under approximate-DP. We propose the use of stochastic gradient (SGD) with a mean estimation gradient oracle based on the results in Section 3. This technique yields nearly-tight bounds in the convex case (similar to first row of Table 2), and for the nonconvex case the stationarity rates are nearly dimension independent (last row of Table 2). The main challenge here is the bias in mean estimation, which dramatically deteriorates the rates of SGD. Hence we propose a bias-reduction method inspired by the simulation literature [9]. This technique uses a random batch size in an exponentially increasing schedule and a telescopic estimator of the gradient which—used in conjunction with our DP mean estimation methods—provides a stochastic first-order oracle that attains bias similar to the one of a full-batch algorithm, with moderately bounded variance. Note that using the full-batch in this case would lead to polynomially weaker rates; in turn, our method leverages the batch randomization to conduct a more careful privacy accounting based on subsampling and the fully-adaptive properties of DP [44]. The introduction of random batch sizes and the random evolution of the privacy budget leads to various challenges into analyzing the performance of SGD. First, we analyze a randomly stopped method, where the stopping time dictated by the privacy budget. Noting that the standard SGD analysis bounds the cumulative regret, which is a submartingale, we carry out this analysis by integrating ideas from submartingales and stopping times [37]. Second, this analysis only yields the desired rates with constant probability. Towards high-probability results, we leverage a private model selection [31] based on multiple runs of randomly-stopped SGD that exponentially boosts the success probability (details in Section 6).
In Section 7, we study further DP-SO and DP-ERM algorithms for the convex case. Our algorithms are based on regularized output perturbation with an $\ell_{\infty}$ projection post-processing step. While this projection step is rather unusual, its role is clear from the analysis: it leverages the $\ell_{\infty}$ bounds of noise addition, which in conjunction with convexity provides an error guarantee that also leverages the gradient sparsity. This algorithm is nearly-optimal for approximate DP. For pure-DP, the previous algorithm requires an additional smoothness assumption, hence we propose a second algorithm based on the exponential mechanism [33] run over a net of suitably sparse vectors. Neither of the pure-DP algorithms matches the lower bound for mean estimation (the gap in the exponent of the rate is of $1/6$), but they attain the first nearly dimension-independent rates for this problem.
@@ 1.2 Related Work
DP optimization is an extensively studied topic for over a decade (see [6, 7, 20], and the references therein). In this field, some works have highlighted the role of model sparsity (e.g., using sparsity-promoting $\ell_{1}$-ball constraints) in near-dimension independent excess-risk rates for DP optimization, both for ERM and SCO [24, 40, 41, 4, 8, 15, 14]. These settings are unrelated to ours, as sparse predictors are typically related to dense gradients.
Another proposed assumption to mitigate the impact of dimension in DP learning is that gradients lie (approximately) in a low dimensional subspace [47, 26, 30, 29] or where dimension is substituted by a bound on the trace of the Hessian of the loss [32]. These useful results are unfortunately not applicable to our setting of interest, as we are interested in arbitrary gradient sparsity patterns for different datapoints.
Substantially less studied is the role of gradient sparsity. Closely related to our work, Zhang et al. [46] studied approximate DP-ERM under gradient sparsity, with some stronger assumptions. Aside from an additional $\ell_{\infty}$ bound on individual gradients, the following partitioning sparsity assumption is imposed: the dataset $S$ can be uniformly partitioned into subsets $S_{1},\ldots,S_{m}$ with a uniform gradient sparsity bound: for all $k\in[m]$ and $x\in{\mathcal{X}}$, $\big{\|}\sum_{z\in S_{k}}\nabla f(x,z)\big{\|}_{0}\leq c_{1}$. The work shows poly-logarithmic in the dimension rates, for both convex and nonconvex settings: for convex objectives the rate is $\frac{\mbox{\footnotesize polylog}(d)}{\varepsilon\sqrt{n}}$, and in the nonconvex case is $\frac{\mbox{\footnotesize polylog}(d)}{\sqrt{\varepsilon}n^{1/4}}$. Our results only assume individual gradient sparsity, so on top of being more general, they are also faster and provably nearly optimal in the convex case. Another relevant work is Ghazi et al. [21], which studies the computational and utility benefits for DP with sparse gradients in neural networks with embedding tables. With the caveat that variable selection on stochastic gradients is performed at the level of contributing buckets (i.e. rows of the embedding table), rather than on gradient coordinates, this work shows substantial improvements on computational efficiency and also on the resulting utility.
@@ 1.3 Future Directions
First, we conjecture that for approximate-DP mean estimation—similarly to the pure-DP case—a lower bound $\Omega\big{(}\sqrt{s\log(d/s)\ln(1/\delta)}/[n\varepsilon]\big{)}$ should exist; such construction could be bootstrapped with a block-diagonal dataset for a tight lower bound (Lemma 4.3). Second, for pure DP-SCO, we believe there should exist an algorithm attaining rates analog to those for mean estimation. Unfortunately, most of variants of output perturbation (including phasing [20, 4, 5]) cannot attain such rates. From a practical perspective, the main open question is whether our rates may be attained without prior knowledge of $s$; note that all our mean estimation algorithms (which carries over to our optimization results) depend crucially on knowledge of this parameter. While we can treat $s$ as a hyperparameter, it would be highly beneficial to design algorithms that automatically adapt to it.
We believe our bias-reduction is of broader interest. For example, [28, 34] have shown strong negative results about bias in DP mean estimation. While similar lower bounds may hold for sparse estimation, bias-reduction allows us to amortize this error within an iterative method, preventing error accumulation.
@@ 3 Upper Bounds for DP Mean Estimation with Sparse Data
We first study DP mean estimation with sparse data. Our first result is that the projection mechanism [35] is nearly optimal, both for pure- and approximate-DP. In our case, we interpret the marginals on each of the $d$ dimensions as the queries of interest: this way, the $\ell_{2}$-error on private query answers corresponds exactly to the $\ell_{2}$-norm estimation error. A key difference to the approach in [35] and related works is that we project the noisy answers onto the set ${\cal K}:={\cal B}_{1}^{d}(0,L\sqrt{s})$, which is a (coarse) convex relaxation of $\mathrm{conv}({\cal S}_{s}^{d})$. This is crucial to make our algorithm efficiently implementable.
@@
return $\hat{z}=\operatorname*{\mathrm{argmin}}\{\|z-\tilde{z}\|_{2}:z\in{\cal K}\}$, where ${\cal K}:={\cal B}_{1}^{d}(0,L\sqrt{s})$
@@ 3.1 Sharper Approximate DP Mean Estimation via Compressed Sensing
We propose Algorithm 2, a more accurate method for approximate-DP mean estimation based on compressed sensing [45]. The precise improvements relate to reducing the $\log(d)$ factor to $\log(d/s)$, and a faster rate dependence on the confidence $\beta$. The idea is that for sufficiently high dimensions a small number of random measurements suffices to estimate a noisy and approximately sparse signal. These properties follow from existing results in compressed sensing, which provide guarantees based on the $\ell_{2}$-norm of the noise, and the best sparse approximation in the $\ell_{2}$-norm (known as $\ell_{2}\text{-}\ell_{2}$-stable and noisy recovery) [45]. We will exploit such robustness in two ways: regarding the noise robustness, this property is used in order to perturb our measurements, which will certify the privacy; on the other hand, the approximate recovery property is used to find a sparser approximation of our empirical mean. As the approximation is only used for analysis, we can resort on the Approximate Caratheodory Theorem to certify the existence of a sparse vector whose sparsity increases more moderately with $n$ than the empirical average [36].
@@
$\bar{z}(S)=\frac{1}{n}\sum_{i\in[n]}z_{i}\in\mathbb{R}^{d}$ from dataset $S\in({\cal S}_{s,d})^{n}$; privacy parameters $\varepsilon,\delta>0$
@@
If $6\exp\{-cm\}\leq\delta<\frac{s\ln(d/s)}{m^{2}}$ (where $c>0$ is a constant) and $0<\varepsilon\leq 1$, then Algorithm 2 is $(\varepsilon,\delta)$-DP, and with probability $1-\delta/2-\beta$,
@@
We now proceed to the accuracy guarantee. By [45, Theorem 3.6 (b)], under the same event as stated above (which has probability $21-\delta/$ 2) we have